# Cross-Domain Few-Shot Learning with Coalescent Projections and Latent Space Reservation

### v1

@@ 4.2.1 Novel Class Generation in Latent-space
Motivated by [27] and [2], we propose generating novel pseudo-classes by mixing the distributions. The purpose of the pseudo novel embeddings is to repel the embeddings of the current source dataset, compact the space they occupy, making room for the real novel classes that the model will encounter in the future by stretching the other areas of the embedding space.
@@
$\mu_{k}$ $:=\frac{1}{N_{k}}\sum_{y_{i}=k}h(x_{i}),$ (5)
$\operatorname{cov}_{k}$ $:=\frac{1}{N_{k}-1}\sum_{y_{i}=k}{(h(x_{i})-\mu_{k})}^{T}(h(x_{i})-\mu_{k}),$ (6)
where :=assign:=:= is the assignment operator, $h(x_{i})$ is the embedding for an $i$-th input sample $x_{i}$, $y_{i}$ is the corresponding label, and $\mu_{k}$ and $\operatorname{cov}_{k}$ are the mean (prototype) and covariance of the class $k$, respectively.
First, we start the process by generating a pool of $G$ pseudo-distributions by calculating a linear combination of pairs of base class distributions as follows:
@@
$\operatorname{cov}_{\tilde{k}}$ $:=\alpha\operatorname{cov}_{a}+(1-\alpha)\operatorname{cov}_{b},$ (8)
where $\tilde{k}$ is the generated pseudo class candidate, $\alpha\sim U(0,1)$ and $U$ is the uniform distribution. Therefore, the distribution of each pseudo-class can be defined as $\{(\mu_{\tilde{k}}^{i},\operatorname{cov}_{\tilde{k}}^{i})\}_{i=1}^{G}$.
We can sample from these distributions to obtain the pseudo-samples $\{(x_{i},y_{i})\}_{i=1}^{GK}$, where $x_{i}$ and $y_{i}\in\{1,2,\dots,G\}$ are the pseudo-embeddings and their corresponding labels, respectively. However, since these pseudo-classes might be similar to each other or to the base classes, they do not have sufficient information gain. Therefore, we introduce two crucial criteria to filter the candidate distributions to achieve more useful embeddings. First, the prototypes should be diverse and distinct from one another to convey more useful information and cover other areas of the latent space. Second, they should not also be similar to the distributions of the base classes. In the following, we introduce the novel-novel and novel-base criteria to diversify the pseudo-novel classes.
To ensure classes are sufficiently distinct from one another, we first assign them similarity scores and then prune the $G$ candidates based on the calculated scores. By having the $P\in\mathbb{R}^{G\times d}$ matrix of pseudo prototype candidates, we can calculate their similarity as follows:
$S$ $:=PP^{T},$ (9)
$S$ $:=S-(S\circ I),$ (10)
where $S$ is a ${G\times G}$ matrix, and $\circ$ denotes the element-wise (Hadamard) product. The purpose of the Eq. 10 is to ignore the self-similarity by zeroing the diagonal elements of the similarity matrix $S$. Next, we calculate the sum of each row of this symmetric matrix with
@@
where $\vec{1}$ is a $G\times 1$ column vector of ones. This score matrix indicates the degree of similarity between each generated pseudo-prototype to the others. Therefore, we choose $N_{0}\times N$ minimum scores to obtain an $N_{0}\times N$-ways episode.
To obtain base-novel scores to further refine the pool of candidates, we calculate the total divergence between the distribution of the pseudo novel class $\tilde{k}$ and all base classes as follows:
$\begin{split}&\operatorname{Div}_{\tilde{k}}:=\sum_{k=1}^{K_{\text{Base}}}D_{\text{KL}}(X_{k}\parallel X_{\tilde{k}})=\frac{1}{2}\sum_{k=1}^{K_{\text{Base}}}\Bigl{[}\\ &\operatorname{Trace}(\operatorname{cov}_{\tilde{k}}^{-1}\operatorname{cov}_{k})-d+\\ &(\mu_{k}-\mu_{\tilde{k}})^{T}\operatorname{cov}_{\tilde{k}}^{-1}(\mu_{k}-\mu_{\tilde{k}}))+\ln\frac{|\operatorname{cov}_{\tilde{k}}|}{|\operatorname{cov}_{k}|}\Bigr{]},\end{split}$ (12)
where $D_{\text{KL}}(.\parallel.)$ indicates the K-L divergence between two distributions, and |.||.|| . | denotes the determinant operator. These scores indicate the degree to which a pseudo-novel distribution differs from all base classes. By choosing $N$ distributions with the highest scores, we will have $N$ distributions out of the pool of $G$ distributions for $N$ classes.
The pseudo-embeddings for each pseudo-class in the one episode can now be generated by sampling from each distribution $K$ and $M$ times for the pseudo-novel support and query sets, respectively.
To facilitate the process, we calculate these pseudo-embeddings for the $E_{\text{P}}$ episodes at the beginning to create a dataset of pseudo-episodes. We include the pseudo-embeddings in each pseudo-episode in the calculation of the prototypical loss alongside the embeddings of the base classes in each episode.
Figure 3: The calculation of the pseudo-embeddings. Circles and stars denote the embeddings and generated pseudo-embeddings, respectively. The bold arrows show the repulsive forces.
@@ 4.2.2 Novel Class Generation in Input Space
To fully utilize the previous components, we need to achieve two additional goals. First, we require more diverse embedding locations to reserve additional parts of the latent space and push away from and improve the decision boundaries around the embeddings of the base dataset. Second, since our network is well-trained on natural images with DINO, the improvements resulting from the deployment of pseudo-classes may not be measurable, as the accuracy is already saturated close to 100% in the base domain. Thus, we should make the classification more challenging for the network.
Inspired by [40] and [29], we utilize a straightforward yet effective SST during meta-training on the base dataset, which satisfies both requirements. Here, we generate three additional rotated images for each sample per class and assign them new labels.
$\{(\operatorname{Rotate}(X_{i},d\times 90^{\circ}),dL+Y_{i})\},$ (13)
[…]
The new rotated photos are mapped to the new areas in the latent space and cover these areas. They reserve the space effectively and impose more nuanced and complicated boundaries for the base classes. Therefore, when the network encounters novel classes, it relies on the unique characteristics of those embeddings compared to those of the observed base classes. Besides, they are difficult to classify for the network, just as they are for humans; hence, deploying the rotated samples can also simulate encountering new images from unseen domains. Therefore, we can measure the small improvement in the validation phase. Furthermore, we do not lose information, unlike when performing random cropping, because we utilize the entire image after the rotations [15]. Ultimately, this method is very well aligned with the prototypical classification. The network is not forced to drastically change the location of the mapped samples, even if we have more samples for each novel class.

### v2

@@ 4.2.1 Novel Class Generation in Latent-Space
Motivated by [verma2019manifold] and [chen2025pseudo], we propose generating novel pseudo-classes by mixing the distributions. The purpose of the pseudo-novel embeddings is to repel the embeddings of the current base dataset, compact the space they occupy, and make room for the real novel classes that the model will encounter in the future by stretching the other areas of the embedding space.
@@
$\mu_{k}$ $:=\frac{1}{N_{k}}\sum_{y_{i}=k}(x_{i}),$ (5)
$\Sigma_{k}$ $:=\frac{1}{N_{k}-1}\sum_{y_{i}=k}{((x_{i})-\mu_{k})}((x_{i})-\mu_{k})^{\text{T}},$ (6)
where :=:= is the assignment/definition operator, $(x_{i})$ is the embedding for an $i$-th input sample $x_{i}$, $y_{i}$ is the corresponding label, and $\mu_{k}$ and $\Sigma_{k}$ are the mean (prototype) and covariance of the class $k$, respectively.
First, we start the process by generating a pool of $N_{\mathrm{C}}$ pseudo-distributions by calculating a linear combination of pairs of base class distributions as follows:
@@
$\Sigma_{\tilde{k}}$ $:=\alpha\Sigma_{a}+(1-\alpha)\Sigma_{b},$ (8)
where $\tilde{k}$ is the generated pseudo class candidate, $\alpha\sim U(0,1)$ and $U$ is the uniform distribution. The distribution of each pseudo-class can be defined as $\{(\mu_{\tilde{k}}^{i},\Sigma_{\tilde{k}}^{i})\}_{i=1}^{N_{\mathrm{C}}}$.
While one may directly sample from these distributions to obtain the pseudo-samples $\{(x_{i},\tilde{k})\sim G_{\tilde{k}}\}_{i=1}^{K}$, where $G_{\tilde{k}}$ is the $\tilde{k}-$th pseudo-distribution (pseudo-class $\tilde{k}$), since these pseudo-classes may be similar to each other or to the base classes, they may not have sufficient information gain. Therefore, we introduce two crucial criteria to filter the candidate distributions to obtain more useful embeddings. First, the prototypes should be diverse and distinct from one another to convey more useful information and cover other areas of the latent space. Second, they should not resemble the distributions of the base classes. In the following, we introduce the novel-novel and novel-base criteria to diversify the pseudo-novel classes.
To ensure classes are sufficiently distinct from one another, we first assign them similarity scores and then prune the $N_{\mathrm{C}}$ candidates based on the calculated scores. By having the $P\in\mathbb{R}^{N_{\mathrm{C}}\times d}$ matrix of pseudo prototype candidates, we can calculate their similarity as follows:
$S$ $:=PP^{\text{T}},$ (9)
$S$ $:=S-(S\odot I),$ (10)
where $S$ is a $N_{\mathrm{C}}\times N_{\mathrm{C}}$} matrix, and $\odot$ denotes the element-wise (Hadamard) product. The purpose of the Eq. 10 is to ignore the self-similarity by zeroing the diagonal elements of the similarity matrix $S$. Next, we calculate the sum of each row of this symmetric matrix with
@@
where $\vec{1}$ is a $N_{\mathrm{C}}\times 1$ column vector of ones. This score matrix indicates the degree of similarity between each generated pseudo-prototype and the others. Therefore, we choose $N_{0}\times N$ minimum scores to obtain an $N_{0}\times N$-ways episode.
To obtain base-novel scores to refine the pool of candidates further, we calculate the total divergence between the distribution of the pseudo-novel class $\tilde{k}$ and all base classes as follows:
$\begin{split}\xi_{\tilde{k}}&:=\sum_{k=1}^{N_{\text{Base}}}D_{\text{KL}}(X_{k}\parallel X_{\tilde{k}})=\frac{1}{2}\sum_{k=1}^{N_{\text{Base}}}\Bigl[\operatorname{Tr}(\Sigma_{\tilde{k}}^{\dagger}\Sigma_{k})\\ &+(\mu_{\tilde{k}}-\mu_{k})^{\text{T}}\Sigma_{\tilde{k}}^{\dagger}(\mu_{\tilde{k}}-\mu_{k}))+\ln\frac{|\Sigma_{\tilde{k}}|}{|\Sigma_{k}|}-d\Bigr],\end{split}$ (12)
where $D_{\text{KL}}(.\parallel.)$ indicates the K-L divergence between two distributions, $\operatorname{Tr}$ is the trace of a matrix, $|.|$ denotes the determinant operator, $N_{\text{Base}}$ is the number of base classes, and $\dagger$ is the pseudoinverse [goodfellow2016deep]. Moreover, we use ridge regression [hoerl1970ridge] by adding $\epsilon I$ to the covariance matrices for numerical stability, where $\epsilon$ is a small scalar. These scores indicate the degree to which a pseudo-novel distribution differs from all base classes. By choosing $N$ distributions with the highest scores, we will have $N$ distributions out of the pool of $N_{\mathrm{C}}$ distributions for $N$ classes.
The pseudo-embeddings for each pseudo-class in the one episode can now be generated by sampling from each distribution $K$ and $K_{q}$ times (shots) for the pseudo-novel support and query sets, respectively. Since the pseudo-embeddings lack corresponding inputs for direct backpropagation, they remain in the same regions and affect the network’s learnable parameters by pushing the embeddings of real input samples away in the loss function. Fig. 3 shows the interactions between the pseudo-embeddings and real embeddings.
Figure 3: Latent Space Reservation. The arrows show the repulsive forces.
@@ 4.2.2 Novel Class Generation in Input Space
To fully utilize the previous components, we need to achieve two additional goals. First, we require more diverse embedding locations to reserve additional parts of the latent space and push away from and improve the decision boundaries around the embeddings of the base dataset. Second, since our network is well-trained on natural images with DINO, the improvements from deploying pseudo-classes may not be measurable, as the accuracy is already close to 100% in the base domain. Thus, we should make the classification more challenging for the network.
Inspired by [zhu2025pass++] and [wang2023rotation], we utilize a straightforward yet effective SST during meta-training on the base dataset, satisfying both requirements. Here, we generate three additional rotated images for each sample per class and assign them new labels.
$\{(\operatorname{Rotate}(x_{i},j\times 90^{\circ}),jL+y_{i})\}_{i=1}^{N_{\text{dataset}}},j\in\{0,1,2,3\},$ (13)
where $L$ is the number of total classes, $N_{\text{dataset}}$ is the number of samples in the dataset, and $y_{i}\in\{0,1,\ldots,L\}$ is the label of the $i\text{-th}$ sample. Therefore, this operation quadruples the number of classes.
The new rotated photos are mapped to the new areas in the latent space and cover these areas. They reserve space effectively and impose more nuanced and complicated boundaries for the base classes. Therefore, when the network encounters novel classes, it relies on the unique characteristics of those embeddings compared to those of the observed base classes. Besides, they are difficult for the network to classify, just as they are for humans; hence, deploying rotated samples can also simulate encountering new images from unseen domains. Therefore, we can measure the small improvement in the validation phase. Furthermore, we do not lose information, unlike in random cropping, because we utilize the entire image after the rotations [lee2020self]. Ultimately, this method is very well aligned with the prototypical classification. The network is not forced to drastically shift the locations of the mapped samples, even when we have more samples per novel class.